OpenStudy (anonymous):

square root 8x - 20 - square root 12 answers: x = –1, x = 15, x = 4, or None of the above

OpenStudy (pfenn1):

Do you mean\[\sqrt{8x-20}=\sqrt{12}\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

\[\sqrt{8x-20}=\sqrt{(2\times2\times2)x-(2\times2\times5)}=2\sqrt{2x-5}\] \[\sqrt{12}=\sqrt{2\times2\times3}=2\sqrt{3}\]

OpenStudy (anonymous):

so which one of the answers i put above is it?

OpenStudy (anonymous):

4?

OpenStudy (pfenn1):

\[\sqrt{8x-20}=\sqrt{12}\] can be simplified to \[2\sqrt{2x-5}=2\sqrt{3}\]

OpenStudy (pfenn1):

Can you simplify it further?

OpenStudy (anonymous):

so...which one is it? no i cant :(

OpenStudy (pfenn1):

subtract 2 from both sides and then square each side and solve for x

OpenStudy (anonymous):

can you show me?

OpenStudy (pfenn1):

if you subtract 2 from both sides you get\[\sqrt{2x-5}=\sqrt{3}\]

OpenStudy (anonymous):

so then it would be -1, or none of the above?

OpenStudy (pfenn1):

square each side to get\[2x-5 = 3\] and solve for x to get

OpenStudy (pfenn1):

\[x=\left( 5+3 \over 2 \right)\]

OpenStudy (anonymous):

6.2?

OpenStudy (pfenn1):

\[x=\left( 3+5\over2 \right)=\left( 8 \over 2 \right)=4\]

OpenStudy (anonymous):

can you help me with another?

OpenStudy (pfenn1):

sure

OpenStudy (anonymous):

Solve. 

OpenStudy (anonymous):

thats a 7 btw

OpenStudy (pfenn1):

\[-5\sqrt{x-9}-7 = 23\] Is this correct?

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

ok \[-5\sqrt{x-9}=30\] divide each side by -5\[\sqrt{x-9}=-6\]

OpenStudy (pfenn1):

can you solve for x from here?